OpenStudy (anonymous):

TRUE or FALSE! The law of cosines reduces to the Pythagorean theorem when the triangle is right. **my answer: FALSE! is that right? @terenzreignz :)

terenzreignz (terenzreignz):

What's the law of cosines? \[\Large c^2 = a^2+b^2 -2ab \ \cos(C)\]

terenzreignz (terenzreignz):

So, if it's a right angle, then we have C = 90 degrees... \[\Large c^2 = a^2+b^2-2ab \ \cos(90^o)\]

OpenStudy (anonymous):

yup haha thats it :P

terenzreignz (terenzreignz):

What's cos(90) = ?

OpenStudy (anonymous):

=0 ?

terenzreignz (terenzreignz):

That's right, so this part just cancels out... \[\Large c^2 = a^2 + b^2 \cancel{-2ab \ \cos(90^o)}\] Leaving you with...?

OpenStudy (anonymous):

pythagorean theorem! :) so this is actually TRUE! ahh i see :)

terenzreignz (terenzreignz):

haha

OpenStudy (anonymous):

pythagorean theorem! :) so this is actually TRUE! ahh i see :)